Total number of electrons which will together weigh for one mole of CO2 

Grade:11

1 Answers

Arun
5 years ago

 Mass of an electron = 9.1 x 10-28 g

Or

9.1 x 10-28 g contains = 1 electron

Therefore 1g contains = 1/9.1 x 10-28 *1 = 1.098 x 1027 electrons

 

Now you can do the reat of calculations